[figure: 9.7 cm]

ELECTRONIC STRUCTURE AND TRANSPORT PROPERTIES

OF UFe_{2} SYSTEM

A. Szlaferek, A. Szajek, J. Baszyński, L. Smardz, A. Kowalczyk, M. Timko^{∗}

Institute of Molecular Physics, Polish Academy of Sciences Smoluchowskiego 17, 60-179 Poznań, Poland ∗Institute of Experimental Physics, Slovak Academy of Sciences, Watsonova 47, 043 53 Košice, Slovakia

The electronic structure of UFe_{2} compound has been studied by X-ray photoemission spectroscopy (XPS) and ab initio self-consistent tight binding linear muffin tin orbital (TB LMTO) method. This compound crystallizes in the cubic Laves phase. UFe_{2} is a ferromagnet with T_{c}= 175 K and saturation magnetization M_{s}= 1.09 μ_{B} at T = 4.2 K [1]. The valence bands are compared with the calculated density of states. The uranium f electrons form a wide band with the centre of gravity lying above the Fermi level. The iron d- and uranium f electrons have the main contribution to the total density of electronic states at the Fermi level (2/3 and 1/3, respectively). Contrary to USi_{3} [2,3] the uranium 6p electrons forms wider band which is located between 17.5 and 20 eV (for USi_{3} 1.5 and 19.4 eV) below the Fermi level. The transport properties will be analyzed based on temperature dependence of resistivity.

[1] M.Timko, A.Szlaferek, A.Zentko, J.Kováč, J.Magn.Magn.Mat., (accepted for publ.)

[2] J.A. Morkowski, A. Szajek, A. Jezierski, Acta Phys. Slov. 46, 109 (1996)

[3] J.A. Morkowski, A. Szajek, Acta Phys. Polon., 46, 109 (1996)

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : A. Szlaferek

Address for correspondence : Institute of Molecular Physics, Polish Academy of Sciences Smoluchowskiego 17, 60-179 Poznan, Poland

Email address : szlafer@ifmpan.poznan.pl